# Best Ergodic Averages via Optimal Graph Filters in Reversible Markov Chains

### v1

@@ VI Filter Design with Legendre Polynomials
Here, the $l_{2}$-norm is with respect to the Lebesgue measure over $[\lambda_{{\mathsf{low}}},2]$:
$\|p\|_{2}^{[\lambda_{{\mathsf{low}}},2]}\coloneqq\left(\int_{[\lambda_{{\mathsf{low}}},2]}|p(x)|^{2}\,m(dx)\right)^{1/2}.$
In general, for any Lebesgue integrable weight function $w:[\lambda_{{\mathsf{low}}},2]\rightarrow(0,\infty)$, we can define the weighted $l_{2}$-norm as
$\|p\|_{2,w}^{[\lambda_{{\mathsf{low}}},2]}\coloneqq\left(\int_{[\lambda_{{\mathsf{low}}},2]}|p(x)|^{2}\,w(x)\,m(dx)\right)^{1/2}.$
Therefore, depending on $w$, the solution of the optimization problem $\mathbf{(\widehat{OPT}_{t})}$ varies. However, because we lack knowledge of the eigenvalues of $P$ in the interval $[\lambda_{{\mathsf{low}}},2]$, we do not know how to assign weights to the points within this interval. Thus, it is more reasonable to assign equal weights, i.e., $w(x)=1$, to all the points in $[\lambda_{{\mathsf{low}}},2]$, which leads to the classical $l_{2}$-norm.
To solve the above optimization problem, we need to define orthonormal polynomials over $[\lambda_{{\mathsf{low}}},2]$ with respect to the classical $l_{2}$-norm. Note that if our interval is $[-1,1]$, then Legendre polynomials $\{L_{n}\}$ are the orthogonal polynomials [19, p. 53], where these polynomials are recursively defined as follows:
$L_{0}(z)=1,\,\,L_{1}(z)=z,\,\,$
$(n+1)L_{n+1}(z)=(2n+1)zL_{n}(z)-nL_{n-1}(z)\,\,n\geq 1.$
Note that for each $n\geq 0$, we have [19, Exercise 2.11, p. 62]
@@
$\hat{L}_{n}(z)\coloneqq\sqrt{n+\frac{1}{2}}\,L_{n}.$
@@
Now, we can transform the orthonormality properties of these polynomials to the interval $[\lambda_{{\mathsf{low}}},2]$ by simply applying a linear transformation, which leads to the following orthonormal (with respect to the classical $l_{2}$-norm) polynomials over $[\lambda_{{\mathsf{low}}},2]$:
$\tilde{L}_{n}(z)\coloneqq\frac{2}{2-\lambda_{{\mathsf{low}}}}\,\hat{L}_{n}\left(\frac{2z-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\right)\,\,n\geq 0.$
@@
$\sqrt{\frac{2(n+1)^{2}}{2n+3}}\tilde{L}_{n+1}(z)$
$=\sqrt{2(2n+1)}\,\frac{2z-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\,\tilde{L}_{n}(z)-\sqrt{\frac{2n^{2}}{2n-1}}\tilde{L}_{n-1}(z).$
@@
$\alpha_{n}\coloneqq\sqrt{\frac{2n^{2}}{2n-1}},\,\,\beta_{n}\coloneqq\sqrt{2(2n+1)}\,\,n\geq 1,$
@@
$\xi_{k}^{t-1}\coloneqq\frac{\tilde{L}_{k}(0)}{\sum_{k=0}^{t-1}\tilde{L}_{k}(0)^{2}}.$
@@
$p_{t-1}^{*}(z)\coloneqq\sum_{k=0}^{t-1}\xi_{k}^{t-1}\,\tilde{L}_{k}(z).$
Theorem 4.
We have $p_{t-1}^{*}\in\mathop{\rm arg\,min}{\bf(\widehat{OPT}_{t})}$ for any $t\geq 2$. Hence, the optimal polynomial filter is given by weighted combination of properly normalized and scaled Legendre polynomials of degree at most $t-1$, at each time step $t\geq 2$.
Proof.
@@
Note that any polynomial $p\in P_{t-1}[\lambda_{{\mathsf{low}}},2]$ can be written
$p=\sum_{k=0}^{t-1}c_{k}\,\tilde{L}_{k}.$
If $p$ is a feasible point of ${\bf(\widehat{OPT}_{t})}$, then
$p(0)=\sum_{k=0}^{t-1}c_{k}\,\tilde{L}_{k}(0)=1.$
Hence, by Cauchy’s inequality, we have
@@
$c_{k}=\frac{\tilde{L}_{k}(0)}{\sum_{k=0}^{t-1}\tilde{L}_{k}(0)^{2}}\eqqcolon\xi_{k}^{t-1}.$
This completes the proof. ∎
@@
$\gamma_{K}\coloneqq\frac{\sum_{k=0}^{K}\tilde{L}_{k}(0)^{2}}{\sum_{k=0}^{K+1}\tilde{L}_{k}(0)^{2}}.$
@@
This formula allows us to compute the output of the optimal filter recursively, greatly reducing the algebraic computation needed in each iteration, as it avoids recomputing $p_{K+1}^{*}(L)f$ from the beginning. By storing $p_{K}^{*}(L)f$, $\tilde{L}_{K}(L)f$, and $\tilde{L}_{K-1}(L)f$ in memory, we can efficiently compute $p_{K+1}^{*}(L)f$ using the coupled recursive formulas.
VII Numerical Experiments
In this section, we numerically test the performance of the optimal filters derived in the previous sections. Our goal is to compare these filters with the iterative process described in the classical ergodic theorem. To achieve this, we consider two specific examples, both of which are detailed in Section II: the random walk on graphs and the Glauber chains.
First, we examine the random walk on graphs. Next, we turn our attention to the Glauber chains. By comparing the outcomes from these two examples, we gain a deeper understanding of the strengths and limitations of the optimal filters. This comparison also provides insights into the practical applications of these filters in various Markov chains.
VII-A Random Walk on a Graph
In this section, we consider a very simple random walk on a cycle [17, Example 2.1]. Let $p$ be an odd number. We place the integers $\mathbb{Z}_{p}\coloneqq\{0,1,2,\ldots,p-1\}$ on a circle in the given order and connect each consecutive point in the circle. This forms our graph. In Figure 1, an example is given where $p=11$.
@@ VII-A Random Walk on a Graph
and $\pi(x)=1/p$. For this Markov chain, by [17, Corollary 1] (see also [17, p. 44]), one can obtain the following lower bound to the second smallest eigenvalue $\lambda_{x_{**}}$ of $L$ as follows:
$\lambda_{x_{**}}\geq\lambda_{{\mathsf{low}}}\coloneqq\frac{8p}{(p-1)^{2}(p+1)}.$
@@
VII-B Glauber Chain
In this section, we consider a Glauber chain on a cycle [21]. Let $p$ be the number of vertices in the cycle. Each vertex in this graph can take two values: $+1$ or $-1$. Hence, the state space of this Markov chain is ${\mathsf{X}}\coloneqq\{+1,-1\}^{p}$. In Figure 3, an instance of a state transition of this Markov chain is given where $p=11$. In these two cycles, only the direction of the top nodes are different, and so, with positive probability, a transition can occur.
@@ VII-B Glauber Chain
Figure 3: A state transition of a Glauber chain on the cycle. Up arrow means $+1$, down arrow means $-1$.
@@
where $H(x)=-\sum_{u,w\in\mathbb{Z}_{p}:u\sim w}J_{v,w}\,x(v)x(w)$ is the energy of the state $x$ and $Z(\beta)\coloneqq\sum_{x\in{\mathsf{X}}}e^{-\beta H(x)}$ is the partition function. Therefore, we have
[…]
where $S(x,w)\coloneqq\sum_{u:u\sim w}J_{u,w}\,x(u)$. For this Markov chain, we can obtain its second largest eigenvalue exactly via computing eigenvalues of some low-dimensional matrix [21, Corollary 6]. Indeed, let $M$ be $p\times p$ matrix whose entries are all zero except immediately above and below the diagonal and non-zero entries have the following description:
$M(i,i-1)$ $=s_{i-1}/(c_{i-1}+c_{i})$
$M(i,i+1)$ $=s_{i}/(c_{i-1}+c_{i}),$
where $J_{i}\coloneqq J_{i,i+1}$, $s_{i}\coloneqq\sinh(2J_{i})$, and $c_{i}\coloneqq\cosh(2J_{i})$. Let the eigenvalues of $M$ be $\gamma_{1}\geq\gamma_{2}\geq\ldots\geq\gamma_{p}$. Then, by [21, Corollary 6], the second largest eigenvalue of $P$ is given by
$1-\frac{1-\gamma_{1}}{p}.$
Hence, we can take $\lambda_{{\mathsf{low}}}$ as
$\lambda_{x_{**}}=\frac{1-\gamma_{1}}{p}\eqqcolon\lambda_{{\mathsf{low}}}.$
@@
VIII Conclusion
In this paper, we explore the optimization of ergodic or Birkhoff averages in the context of the ergodic theorem for reversible Markov chains via utilizing graph signal processing techniques to achieve faster convergence to desired values. We use the transition probabilities of a Markov chain to form a directed graph so that we can represent a function on the state space as a graph signal. This allowed us to introduce a graph variation concept and define the graph Fourier transform for graph signals on this directed graph. We identify the inherent non-optimality of the standard graph filter used in ergodic theorem iterations and address this by formulating three optimization problems with distinct objectives. The solutions to these problems yield the Bernstein, Chebyshev, and Legendre polynomial filters, each offering improved performance over traditional method. Our numerical experiments demonstrate that the Bernstein filter provides a modest improvement over the ergodic average. More significantly, the Chebyshev and Legendre filters show substantial enhancements, achieving rapid convergence to the desired value. These results highlight the potential of using optimized graph filters to improve the efficiency of ergodic averages.
The success of the proposed filters in this study shows the power of graph signal processing in addressing complex problems in stochastic processes and dynamical systems. Future work could extend this approach to other types of Markov chains and more abstract state spaces. Additionally, investigating the practical applications of these optimized filters in various domains, such as statistical physics, algorithm design, and network theory, could provide valuable insights and benefits.
IX Acknowledgements
The author is grateful to Professor Alexander Goncharov for directing him to significant results in approximation theory.

### v2

@@ 6 Filter Design with Legendre Polynomials
Here, the $l_{2}$-norm is with respect to the Lebesgue measure over. In general, for any Lebesgue integrable weight function $w:[\lambda_{{\mathsf{low}}},2]\rightarrow(0,\infty)$, we can define the weighted $l_{2}$-norm as
$\|p\|_{2,w}^{[\lambda_{{\mathsf{low}}},2]}\triangleq\left(\int_{[\lambda_{{\mathsf{low}}},2]}|p(x)|^{2}\,w(x)\,m(dx)\right)^{1/2}.$
Therefore, depending on $w$, the solution of the optimization problem $\mathbf{(\widehat{OPT}_{t})}$ varies. However, because we lack knowledge of the eigenvalues of $P$ in the interval $\lambda_{{\mathsf{low}}},2]$, we do not know how to assign weights to the points within this interval. Thus, it is more reasonable to assign equal weights, i.e., $w(x)=1$, to all the points in $\lambda_{{\mathsf{low}}},2]$, which leads to the classical $l_{2}$-norm.
To solve the above optimization problem, we need to define orthonormal polynomials over $\lambda_{{\mathsf{low}}},2]$ with respect to the classical $l_{2}$-norm. Note that if our interval is [1,1][-1,1], then Legendre polynomials $\{L_{n}\}$ are the orthogonal polynomials [15, p. 53], where these polynomials are recursively defined as follows:
$L_{0}(z)=1,\,\,L_{1}(z)=z,\,\,(n+1)L_{n+1}(z)=(2n+1)zL_{n}(z)-nL_{n-1}(z)\,\,n\geq 1.$
Note that for each $n\geq 0$, we have [15, Exercise 2.11, p. 62]
@@
$\hat{L}_{n}(z)\triangleq\sqrt{n+\frac{1}{2}}\,L_{n}.$
@@
Now, we can transform the orthonormality properties of these polynomials to the interval $\lambda_{{\mathsf{low}}},2]$ by simply applying a linear transformation, which leads to the following orthonormal (with respect to the classical $l_{2}$-norm) polynomials over $\lambda_{{\mathsf{low}}},2]$:
$\tilde{L}_{n}(z)\triangleq\frac{2}{2-\lambda_{{\mathsf{low}}}}\,\hat{L}_{n}\left(\frac{2z-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\right)\,\,n\geq 0.$
@@
$\sqrt{\frac{2(n+1)^{2}}{2n+3}}\tilde{L}_{n+1}(z)=\sqrt{2(2n+1)}\,\frac{2z-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\,\tilde{L}_{n}(z)-\sqrt{\frac{2n^{2}}{2n-1}}\tilde{L}_{n-1}(z).$
@@
$\alpha_{n}\triangleq\sqrt{\frac{2n^{2}}{2n-1}},\,\,\beta_{n}\triangleq\sqrt{2(2n+1)}\,\,n\geq 1,$
@@
$\xi_{k}^{t-1}\triangleq\frac{\tilde{L}_{k}(0)}{\sum_{k=0}^{t-1}\tilde{L}_{k}(0)^{2}}.$
@@
$p_{t-1}^{*}(z)\triangleq\sum_{k=0}^{t-1}\xi_{k}^{t-1}\,\tilde{L}_{k}(z).$
Theorem 6.1.
We have $p_{t-1}^{*}\in\mathop{\rm arg\,min}{\bf(\widehat{OPT}_{t})}$ for any $t\geq 2$. Hence, the optimal polynomial filter is given by weighted combination of properly normalized and scaled Legendre polynomials of degree at most t1t-1, at each time step $t\geq 2$.
Proof 6.2.
@@
Note that any polynomial $p\in P_{t-1}[\lambda_{{\mathsf{low}}},2]$ can be written $p=\sum_{k=0}^{t-1}c_{k}\,\tilde{L}_{k}.$ If $p$ is a feasible point of $\bf(\widehat{OPT}_{t})$}, then $p(0)=\sum_{k=0}^{t-1}c_{k}\,\tilde{L}_{k}(0)=1.$ Hence, by Cauchy’s inequality, we have
@@
$c_{k}=\frac{\tilde{L}_{k}(0)}{\sum_{k=0}^{t-1}\tilde{L}_{k}(0)^{2}}\triangleq\xi_{k}^{t-1}.$
This completes the proof.
@@
$\gamma_{K}\triangleq\frac{\sum_{k=0}^{K}\tilde{L}_{k}(0)^{2}}{\sum_{k=0}^{K+1}\tilde{L}_{k}(0)^{2}}.$
@@
This formula allows us to compute the output of the optimal filter recursively, reducing the algebraic computation needed in each iteration, as it avoids recomputing $p_{K+1}^{*}(L)f$ from the beginning. By storing $p_{K}^{*}(L)f$, $\tilde{L}_{K}(L)f$, and $\tilde{L}_{K-1}(L)f$ in memory, we can efficiently compute $p_{K+1}^{*}(L)f$ using the coupled recursive formulas.
7 Numerical Experiments
In this section, we present numerical experiments to illustrate the performance of the optimal filters derived in the previous sections. Our primary aim is to compare these filters with the standard iterative procedure from the classical ergodic theorem. To this end, we consider two simple yet representative examples: the random walk on graphs and the Glauber chain. These examples are chosen precisely because they capture the essential features of our theoretical framework while remaining transparent and easily interpretable. As our goal is to demonstrate the effectiveness of the method in a clear and accessible manner, we deliberately keep the setup simple. Extending these techniques to larger-scale and real-world applications is a natural next step, and we leave such implementations to practitioners who may build on the foundations laid in this work.
7.1 Random Walk on a Graph
Since the graph is connected, the corresponding Markov chain is irreducible. It is also straightforward to see that the Markov chain is reversible with respect to $\pi(x)=\frac{d(x)}{2|E|},\,\,x\in{\mathsf{X}},$ where $|E|$ is the number of edges.
In the numerical experiments, we consider a very simple random walk on a cycle [6, Example 2.1]. Let $p$ be an odd number. We place the integers $\mathbb{Z}_{p}\triangleq\{0,1,2,\ldots,p-1\}$ on a circle in the given order and connect each consecutive point in the circle. This forms our graph. In Figure 3, an example is given where $p=11$.
@@ 7.1 Random Walk on a Graph
and $\pi(x)=1/p$. For this Markov chain, by [6, Corollary 1] (see also [6, p. 44]), one can obtain the following lower bound to the second smallest eigenvalue $\lambda_{x_{**}}$ of $L$ as follows:
$\lambda_{x_{**}}\geq\lambda_{{\mathsf{low}}}\triangleq\frac{8p}{(p-1)^{2}(p+1)}.$
@@
7.2 Glauber Chain
Let $V$ and $S$ be finite sets. We can think of $V$ as the set of vertices of a graph and $S$ is the set of possible values that can be taken by these vertices. Our state space $\mathsf{X}$} is a subset of $V^{S}$. Let $\pi$ be a probability distribution on $\mathsf{X}$}. In statistical physics, $\pi$ is called a Gibbs distribution.
Here, ${\mathsf{X}}(x,v)\triangleq\{y\in{\mathsf{X}}:y(w)=x(w)\,\,\forall\,w\neq v\}$. Verbally, a new state is chosen according to the probability distribution $\pi$ conditioned on the set of states equal to the current state $x$ at all vertices different from $v$. Define ${\mathsf{X}}(x)\triangleq\bigcup_{v\in V}{\mathsf{X}}(x,v)$. Then, the transition probability can be defined as follows:
In the numerical experiment, we consider a Glauber chain on a cycle [13]. Let $p$ be the number of vertices in the cycle. Each vertex in this graph can take two values: +1+1 or 1-1. Hence, the state space of this Markov chain is ${\mathsf{X}}\triangleq\{+1,-1\}^{p}$. In Figure 5, an instance of a state transition of this Markov chain is given where $p=11$. In these two cycles, only the direction of the top nodes are different, and so, with positive probability, a transition can occur.
@@ 7.2 Glauber Chain
Figure 5: A state transition of a Glauber chain on the cycle. Up arrow means +1+1, down arrow means 1-1.
@@
where $H(x)=-\sum_{u,w\in\mathbb{Z}_{p}:u\sim w}J_{v,w}\,x(v)x(w)$ is the energy of the state $x$ and $Z(\beta)$ is the partition function. Therefore, we have
$P(x,y)=\frac{1}{p}\sum_{w\in\mathbb{Z}_{p}}\frac{e^{\beta y(w)S(x,w)}\,1_{\{x(v)=y(v)\,\,\forall v\neq w\}}}{e^{\beta y(w)S(x,w)}+e^{-\beta y(w)S(x,w)}},$
where $S(x,w)\triangleq\sum_{u:u\sim w}J_{u,w}\,x(u)$. For this Markov chain, we can obtain its second largest eigenvalue exactly via computing eigenvalues of some low-dimensional matrix [13, Corollary 6]. Indeed, let $M$ be $p\times p$ matrix whose entries are all zero except immediately above and below the diagonal and non-zero entries have the following description:
$M(i,i-1)=s_{i-1}/(c_{i-1}+c_{i}),\,\,M(i,i+1)=s_{i}/(c_{i-1}+c_{i}),$
where $J_{i}\triangleq J_{i,i+1}$, $s_{i}\triangleq\sinh(2J_{i})$, and $c_{i}\triangleq\cosh(2J_{i})$. Let the eigenvalues of $M$ be $\gamma_{1}\geq\gamma_{2}\geq\ldots\geq\gamma_{p}$. Then, by [13, Corollary 6], the second largest eigenvalue of $P$ is given by $1-\frac{1-\gamma_{1}}{p}.$ Hence, we can take $\lambda_{{\mathsf{low}}}$ as $\lambda_{x_{**}}=\frac{1-\gamma_{1}}{p}\triangleq\lambda_{{\mathsf{low}}}.$
@@
8 Conclusion
In this paper, we explore the optimization of ergodic or Birkhoff averages in the context of the ergodic theorem for reversible Markov chains via utilizing graph signal processing techniques to achieve faster convergence to desired values. We use the transition probabilities of a Markov chain to form a directed graph so that we can represent a function on the state space as a graph signal. This allowed us to introduce a graph variation concept and define the graph Fourier transform for graph signals on this directed graph. We identify the inherent non-optimality of the standard graph filter used in ergodic theorem iterations and address this by formulating three optimization problems with distinct objectives. The solutions to these problems yield the Bernstein, Chebyshev, and Legendre polynomial filters, each offering improved performance over traditional method. Our numerical experiments demonstrate that the Bernstein filter provides a modest improvement over the ergodic average. More significantly, the Chebyshev and Legendre filters show substantial enhancements, achieving rapid convergence to the desired value.
Future Research Directions
We have identified several promising research directions to extend the work presented in this paper. A primary area of interest involves the analysis of non-reversible Markov chains. In this setting, the graph Laplacian is no longer symmetric, causing its spectrum to move into the complex plane. This shift presents a significant mathematical challenge: the design of optimal graph filters would necessitate the use of complex polynomials. Consequently, the desirable properties and optimality guarantees associated with Bernstein, Chebyshev, and Legendre polynomials, which are central to our current framework, may no longer hold. Pursuing this direction requires a deeper exploration into the approximation theory of complex polynomials.
A further research direction involves extending this problem to abstract state spaces. In this setting, the classical graph structure no longer exists; however, analogous to deterministic dynamical systems, spectral filtering methods may still be applied to accelerate ergodic averages. To adapt the optimal filtering techniques from our graph signal processing framework to this continuous or infinite context, a key challenge arises: the spectral domain, which may contain countably many elements (or even a continuous spectrum), must be approximated by a finite set. This approximation must satisfy a specific spectral gap property to ensure that the filters remain effective. Bridging this gap between discrete graph signals and abstract operator theory would allow for a more generalized application of optimal filtering.
A final research direction involves the investigation of Infinite Impulse Response (IIR) filters, which correspond to rational filtering methods. In this framework, the acceleration is achieved using rational functions rather than polynomials. Implementing this approach requires a deep dive into the approximation properties of rational functions, analogous to the well-studied roles of Bernstein, Chebyshev, and Legendre polynomials in the FIR (Finite Impulse Response) setting. Successfully extending the theory to rational filters would allow for a rigorous performance comparison, enabling us to identify the specific trade-offs, advantages, and limitations of rational versus polynomial filtering for ergodic acceleration.
9 Acknowledgements
The author thanks Professor Alexander Goncharov for guiding him to important findings in approximation theory.